calculate the mass of 1 atom of chlorine. (m cl) not the molecular weigth
Answer #1

Answer:- one atom of Cl = 5.89 * 10-23 g

since we know that

1 mole of Cl = 6.023 * 1023 atoms of Cl = 35.453 g of chlorine

so for one atom of Cl = 35.453g / 6.023 * 1023

one atom of Cl = 5.89 * 10-23 g

if you want answer in amu then

1 atom of Cl = 35.453 amu

please rate the answer

and if you have doubt then feel free to ask